OpenStudy (anonymous):

Find the y coordinates for (-8, ), (-4, ), (4, ) y=3/4x-3 i just need to know what to do when i plug in the x values

OpenStudy (anonymous):

plug in x = -8 in the equation and find out the value of y. post your answer and steps here,.

OpenStudy (anonymous):

y=-3/4(-8)-3 y=(-3/4)(-8/1)-3 am i on the right page now so far

OpenStudy (anonymous):

wait. is the equation y = 3/4x -3 or y = - 3/4x -3 ?

OpenStudy (anonymous):

its -3/4 sorry

OpenStudy (anonymous):

okay, then you are on the right page so far. what is the value of y?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

how did you get that? show me the steps please.

OpenStudy (anonymous):

i found the common denominator?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[y = (\frac{-3}{4} \times -8 ) - 3\]

OpenStudy (anonymous):

y = (24/4) - 3 y = 6-3 = 3

OpenStudy (anonymous):

-3 times -8 = 24

OpenStudy (anonymous):

i tought u put -8 for x

OpenStudy (anonymous):

I did put x = -8

OpenStudy (anonymous):

\[y = \frac{3}{4}x-3\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

with a negative sign on 3/4

OpenStudy (anonymous):

sorry \[y = \frac{-3}{4}x -3\]

OpenStudy (anonymous):

okay, then substitute x = -8 \[y = (\frac{-3}{4} \times -8 ) - 3 \rightarrow y = \frac{24}{4}-3 \rightarrow y = 6-3 = 3\]

OpenStudy (anonymous):

ok give me one second trying to work it on my own

OpenStudy (anonymous):

3/4 x means three-fourths of x. three fourths of 8 is 6

OpenStudy (anonymous):

the second problem is y= -6

OpenStudy (anonymous):

i worked it the same wasy you did the other one

OpenStudy (anonymous):

no, it is 0. how did you get -6 for x = -4

OpenStudy (anonymous):

whats the last one suppose to be

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

ok got it